# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v2

@@ 4.2 Context-Aware Intent Clustering
The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are determined based on the labeled intent clusters. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using K-means or a context-free approach, and these results are divided based on intent labels, forming two clusters with distinct roles. In the second step, the sentences corresponding to each role are clustered again to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent.
@@ 5.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset will be open-sourced upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., “inquire-insurance”), while 622 are categorized as out-of-domain (e.g., “provide-location”), representing general queries that might commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1, which presents several new challenges such as lexical sparsity in short-text clustering, dependence on prior knowledge for context enrichment, and balancing accuracy with computational efficiency. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset.
@@
Dataset examples intents diversity
BANKING77 13,083 77 0.209
@@
Table 1: Comparison of our proposed dataset with existing benchmark intent clustering datasets.
@@ 5.2 Metrics
Category Method NMI #Good
Base models K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
Data-centric ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
Context-Free LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Context-Aware LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
@@ 6.1 LLM Utilities
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
Coherence Evaluation.
To enhance evaluation performance while preserving data privacy, four Chinese open-sourced LLMs Liu et al. (2024) are fine-tuned using LoRA Hu et al. (2022) on a training dataset comprising 1,772 intent clusters annotated as either “good” or “bad”. Table 3 presents performance of fine-tuned LLMs based on 480 unseen clusters. The results suggest that mainstream open-source LLMs can serve as robust evaluators for assessing the quality of intent clusters.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7 chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.
Cluster Naming.
@@
Similar to coherence evaluation, we manually annotated 2,500 clusters with corresponding intent labels to fine-tune and assess the utility of LLMs for cluster naming. Since labels are not unique, the accuracy is manually assessed by four human experts based on the alignment with the true labels in the dataset. The results in Table 4 suggest that fine-tuned LLMs have promising performance in providing intuitive names and aligning closely with human perceptions.
6.2 Intent Clustering Performance
Implementation.
The qwen14B model was utilized for both coherence evaluation and cluster naming. For all clustering methods, semantic-rich embeddings are obtained using the BAAI General Embeddings (BGE) model [^2] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). Among the clustering algorithms, hierarchical clustering demonstrated the highest NMI and goodness scores, thus serving as the base model for the proposed methods. Sampling techniques (i.e., random and convex) are implemented to select representative sentences from each cluster as input for LLM.
@@ 6.2 Intent Clustering Performance
Among data-centric baselines, refining sentence embeddings through keyphrase expansion demonstrated the most improvement. However, deriving pairwise constraints for fine-tuning the embedding model, as proposed in ClusterLLM, proved ineffective due to the excessive number of clusters in the dataset. Interestingly, the evaluation of goodness does not always align with the NMI score, offering additional insights into cluster quality. For example, while keyphrase expansion improved the NMI score, the generated keyphrases could distort the meaning of the original sentence, resulting in less representative clusters and a slightly lower goodness score compared to the base model.
The proposed LLM-in-the-loop intent clustering, abbreviated as the “LLM-ITL” approach, demonstrated satisfactory improvement in NMI and significant enhancement in the goodness score. In a context-free setting, the integration of LLM evaluation and naming outperformed both ClusterLLM and IDAS baselines. By incorporating effective keyphrase expansion, the NMI improved further, highlighting the potential of integrating data-centric methods into the pipeline. Notably, the proposed method does not require prior knowledge of the number of clusters and conducts parameter searches during the clustering process, whereas the baseline methods preset the number of clusters at $n=1507$. This adaptability underscores its significant application value for extracting intent clusters from any noisy text corpora Akama et al. (2020).
Furthermore, context-aware approaches were evaluated with the goal of enhancing the task performance of customer service dialogue intent clustering. A consistent improvement was observed when incorporating dialogue role into the clustering process, resulting in a 4.48% increase in NMI without compromising the goodness score. This highlights the importance of task-centric design in effectively addressing industrial problems.
@@ 6.3 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the baseline method, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
@@ 6.4 Ablation Studies
We present two ablation studies to derive best practices for industrial deployment using different sampling and hyperparameter tuning strategies.
6.4.1 Effective LLM Evaluation
Optimizing the input to LLMs is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 2, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ (for repeated evaluation) or the dimension of convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by various fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 2) when using convex sampling.
6.4.2 Parameter Search Space Pruning
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 10). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 5), indicating a high likelihood that the optimal cluster number is among these candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
@@ 7 Conclusion
This paper tackles the dialogue intent clustering task by leveraging intermediate results through a human-aligned LLM-in-the-loop approach. Extensive experiments show that fine-tuned LLMs perform robustly in semantic coherence evaluations and cluster labeling. Both context-free and context-aware methods surpass existing LLM-guided clustering baselines, significantly improving cluster quality and downstream application performance. The ablation studies further provide practical insights for efficient industrial deployment. Future research is encouraged to enhance LLM reliability by integrating confidence estimation into the evaluation process, as well as employing advanced base algorithms and more sophisticated judgment formats to reveal additional insights.
@@ Appendix A Data Collection and Annotation
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 120,000 calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
@@
To create the intent clustering dataset, we recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
@@
Appendix B Prompt Template and More Results for LLM Utilities
The prompt for coherence evaluation and cluster naming are translated into English for demonstration. For each input, a few-shot demonstration with five input-output pairs is provided to ensure the consistency of the output format and enhance understanding of the task.

### v3

@@ 3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are naturally determined based on the LLM- intent label. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\notin\{\text{"answer-"},\dots\}\}.\end{array}$
In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$, to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent. Finally, the two groups of clusters are merged together, serving as the final or intermediate result to be further processed.
@@ 4.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset and fine-tuned models will be released upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., inquire-insurance), while 622 are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1. It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset and motivates the development of semantic-guided techniques.
@@
Dataset sentences intents diversity
BANK77 3080 77 0.209
@@
Table 1: Comparison of proposed dataset with existing intent clustering datasets.
Category Method NMI #Good
Baselines K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
LLM-Guided Clustering Baselines ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
LLMEdgeRefine Feng et al. (2024) 0.7411 87.2%
Proposed Method (context-free) LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Proposed Method (context-aware) LLM-ITL + merge 0.8420 97.8%
LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.
@@ 4.2 Metrics
4.3 Implementations
Four open-sourced Chinese LLMs [^2] The models are available at: Qwen2.5-7B; Qwen2.5-14B; Baichuan2-7B ; ChatGLM3-6b are fine-tuned using LoRA Hu et al. (2022) on 4 $\times$ Nvidia A100 GPUs. For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used. For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with intent labels. Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering. To reduce computational load, both random and convex sampling methods are implemented to select 20 representative sentences per cluster as input for the LLM.
@@ 5.1 LLM Utilities
Table 3 presents the performance of the fine-tuned LLM coherence evaluator tested on 480 unseen clusters. The results indicate that mainstream open-source LLMs can effectively serve as robust evaluators for assessing the quality of intent clusters and providing human-aligned judgments. For cluster naming, since labels are not unique, accuracy is manually evaluated by four human experts based on alignment with the true labels in the dataset. The results in Table 4 demonstrate that the fine-tuned LLMs show promising performance in generating intuitive names and adhering strictly to the predefined “Action-Objective” format.
5.2 Main Results
@@
Table 2 presents the main results evaluated using the proposed dataset. Among the LLM-guided clustering baselines, the data-centric Keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al. (2024). However, deriving pairwise constraints for fine-tuning the embedding model, as in ClusterLLM, was ineffective due to the excessive number of clusters in the dataset. Interestingly, goodness evaluation does not always align with the NMI score. For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model. This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
The proposed LLM-in-the-loop intent clustering demonstrated satisfactory improvements in NMI and significant enhancements in the goodness score. In a context-free setting, iterative clustering with convex sentence sampling outperformed both ClusterLLM and IDAS baselines. Effective keyphrase data augmentation further improved NMI, highlighting the potential of integrating data-centric methods into the pipeline. Furthermore, the context-aware approaches consistently improved performance by incorporating dialogue roles and cluster merging, resulting in a 4.48% increase in NMI without compromising the goodness score. This emphasizes the importance of task-centric design in leveraging intent labels effectively for the problem of intent clustering.
Notably, the proposed method does not require prior knowledge of the number of clusters and performs parameter searches during clustering, unlike baseline methods that rely on the true number of clusters. This adaptability underscores its strong application potential for extracting intent clusters from noisy text corpora Akama et al. (2020).
Method Bank77 CLINC(I) MTOP(I) Massive(I)
SCCL 81.77 92.94 73.52 73.90
Self-supervise 80.75 93.88 72.50 72.88
ClusterLLM 85.07 94.00 73.83 77.64
IDAS 82.84 92.35 72.31 75.74
LLMEdgeRefine - 94.86 72.92 76.66
ours 82.32 94.12 72.45 78.12
Table 5: Performance (NMI) on English benchmarks.
5.3 Evaluation on English Benchmarks
To validate the effectiveness of the proposed LLM-in-the-loop technique, we evaluate the context-aware (best-performing) method on four widely used English benchmarks FitzGerald et al. (2023). The Llama-7b model is fine-tuned with 800 intent clusters (200 samples from each dataset) annotated by human experts to derive Action-Objective intent labels and bad clusters through perturbation. As shown in Table 5, our method delivers performance comparable to state-of-the-art LLM-guided methods and outperforms on Massive(I) dataset. Furthermore, the computational cost is lower than the compared methods as measured in Section 5.5.
@@ 5.4 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the base model, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
5.5 Analysis of Computational Cost
We systematically compare the cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, with a test set of $S=3,080$ sentences and true cluster number $N=77$. The proposed LLM-in-the-loop iterative method (Algorithm 1) uses candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ LLM calls per iteration, and with $T=3$ iterations, totals $T\cdot\sum_{n_{i}\in N}n_{i}=480$ calls, each with 20 sentences input for coherence evaluation. ClusterLLM, employing a fixed triplet sampling approach with $Q=1,024$, incurs a constant $1,024$ LLM calls. Keyphrase expansion generates one keyphrase per sentence, resulting in a linear cost of $3,080$ LLM calls for the entire dataset. While our method requires more input tokens per LLM call, the cost remains efficient compared to its counterparts since the output token (i.e., “good/bad” label and intent label) is kept concise with fewer LLM calls.
@@ 5.6 Ablation Studies
Effective Coherence Evaluation with Data Sampling and LLM Crowdsourcing
Optimizing the input to LLM evaluator is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ for repeated evaluation with majority voting or the dimension of the convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by four fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 62.50% 65.00% 58.50% 60.00%
Table 6: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters using numerical 1–4 scale scoring.
Binary Judgment vs. Numerical Scoring
To identify the optimal evaluation strategy for our clustering framework, we compared a numerical 1–4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment. We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset. The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting. The 1–4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human evaluation. In contrast, the simpler binary good/bad judgment provided consistent and interpretable quality assessments, proving to be the more effective evaluation strategy.
Method NMI Goodness
Geodesic distance (probabilistic) 0.8420 97.8%
Cosine similarity (deterministic) 0.8102 94.7%
Cosine similarity (probabilistic) 0.8242 95.2%
Geodesic distance (deterministic) 0.8309 95.4%
Table 7: Ablation study on cluster merging methods.
Effectiveness of Cluster Merging Techniques
Table 7 evaluates the impact of distance measures and merging decisions on the proposed cluster merging technique. Results highlight the superiority of geodesic distance in hyperspherical space, demonstrating its ability to effectively capture deeper semantic relationships compared to cosine similarity. Additionally, the probabilistic merging strategy consistently improves robustness by accounting for uncertainty, yielding better clustering quality regardless of the distance measure used. These findings emphasize the importance of both an appropriate distance metric and a probabilistic approach in achieving optimal merging outcomes.
@@
Parameter Search Space Pruning
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 13). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
@@ 6 Conclusion
This paper tackles dialogue intent clustering using a human-aligned LLM-in-the-loop approach. Experiments on a complex, realistic Chinese customer service dataset reveal that fine-tuned LLM utilities excel in semantic coherence evaluation and cluster labeling, with the proposed methods consistently outperforming existing LLM-guided clustering baselines in cluster quality and computational efficiency. Future work should focus on refining evaluation strategies to enhance semantic guidance from LLMs and exploring task-centric approaches for deeper LLM-in-the-loop integration.
Limitations
This work has two main limitations. First, while the coherence evaluation is practical and intuitive, relying solely on cluster coherence as the quality indicator overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important in assessing a cluster’s usefulness. For instance, a cluster labeled “express-feeling” may be overly general and should be refined into more specific categories like “express-appreciation” and “state-dissatisfaction” to enhance interpretability and practical applicability. Additionally, the binary evaluation and numerical scoring considered in this paper are deterministic indicators, overlooking probabilistic judgments and confidence intervals, which could provide deeper insights into clustering assessment and offer greater flexibility for integration into the LLM-in-the-loop clustering process.
Second, the dataset proposed in this study is limited to the Chinese language and focuses on specific domains, namely banking, insurance, and telecommunication. While it represents the largest realistic customer service call dataset for the intent clustering problem, its coverage is restricted in terms of multi-domain and multilingual aspects, potentially affecting the generalizability of our findings to other languages and domains. Furthermore, the scope of this paper is confined to intent clustering, which serves as the initial step in the broader process of intent discovery. Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, are not comprehensively discussed. These areas warrant further investigation to fully reveal the practical values of LLM-in-the-loop design in intent mining tasks.
@@ Appendix A Data Collection and Annotation
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 11,879 customer service calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
@@
To construct the intent clustering dataset, we internally recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
@@
Appendix B Prompt Template
The prompts for coherence evaluation and cluster naming are translated into English for demonstration purposes. Each input is accompanied by a few-shot demonstration with five input-output pairs to ensure consistency in the output format and enhance understanding of the task.